OpenStudy (anonymous):

Convert the decimal to a fraction. Be sure to show all of your work. 0.25

OpenStudy (anonymous):

0.25 = 25/100 = 25%

OpenStudy (anonymous):

.25 is .25 out of a hundred so that can be changed to 25/100 which reads 25 out of one hundred. Reduce to 1/4

OpenStudy (anonymous):

0.25 = 25/100 = 1/4

OpenStudy (anonymous):

ok thanx

OpenStudy (anonymous):

.25 is 25 out of 100....

OpenStudy (anonymous):

hairkat pls assist with my others

OpenStudy (anonymous):

which ones???